If 15 mL of alcohol is added to 50 mL of water,what is the concentration of the solution?

1
Expert's answer
2021-11-08T08:22:00-0500

Molar mass of ethanol = 46.07 g/mol

= 15/46.07= 0.33mol

= 0.33/0.05

= 6.6mol/L